OpenStudy (anonymous):

Tetsuo's savings account has a balance of $1192. After 22 years, what will the amount of interest be at 1.5% compounded annually?

OpenStudy (anonymous):

This is an exponential growth problem. Use the formula A = p(1 + r)^t, where r = rate of interest compounded annually, t = # of years, p = starting amount you put into the account, and A is the total amount of money you have after t years. In your problem, A = 1192(1 + .015)^22...use your calculator to compute that.

OpenStudy (anonymous):

i put it in but its not one of my answers

OpenStudy (anonymous):

Evaluate it correctly on your calculator. Try again.

OpenStudy (anonymous):

The amount of interest is found by taking the total, A, and subtracting 1192 from it, as 1192 is the original amount of money that was put into the account.

OpenStudy (anonymous):

i got it thanks

OpenStudy (anonymous):

Welcome.